# From small eigenvalues to large cuts, and Chowla’s cosine problem

### v1

@@ 1.2 Maximum Cut
There exists $\varepsilon>0$ such that the following holds for every sufficiently large $n$. If $G$ is an $n$-vertex $m$-edge graph with no cut of size larger than $\frac{m}{2}+n^{1+\varepsilon}$, then $G$ is $n^{-\varepsilon}$-close to the disjoint union of cliques.
There is a close relationship between the MaxCut of a graph $G$ and its smallest eigenvalue. It is well known that $\operatorname{surp}(G)\leq|\lambda_{n}|n$ (see e.g. Claim 5.1 for a short proof). For many algebraically defined graph families, we also have $\operatorname{surp}(G)=\Theta(|\lambda_{n}|n)$, but in general these quantities can be far apart. A good way to think about the surplus as a robust version of the smallest eigenvalue: in many natural cases $\operatorname{surp}(G)=\Theta(|\lambda_{n}|n)$, but $\operatorname{surp}(G)$ is much less sensitive to local modifications. Therefore, it is not unexpected that we obtain similar results as in the smallest eigenvalue case, as we present in the next section.
The proof of Theorem 1.3 is presented in Section 12, and the proof of Theorem 1.4 in Section 13.2.
1.3 Smallest eigenvalue
A central topic of spectral graph theory is understanding the structure of graphs, whose adjacency matrix has large smallest eigenvalue. Let $G$ be an $n$-vertex graph and let $\lambda_{n}$ denote the smallest eigenvalue of its adjacency matrix. A simple consequence of the Cauchy-interlacing theorem is that if $G$ is non-empty, then $\lambda_{n}\leq-1$ with equality if and only if $G$ is the disjoint union of cliques. In the 1970’s, Cameron, Goethels, Seidel, and Shult [17] gave a complete characterization of graphs satisfying $|\lambda_{n}|\leq 2$, which are exactly generalized line graphs and some sporadic examples with at most 36 vertices. More recently, Koolen, Yang and Yang [58] obtained a partial characterization in the case $|\lambda_{n}|\leq 3$ by integral lattices. Beyond these specific values, much less is known. Kim, Koolen, and Yang [52] proved the following structure theorem for regular graphs satisfying $|\lambda_{n}|\leq\lambda$. One can find dense induced subgraphs $Q_{1},\dots,Q_{c}$ in $G$ such that each vertex lies in at most $\lambda$ of $Q_{1},\dots,Q_{c}$, and almost all edges are covered by the union of $Q_{1},\dots,Q_{c}$. However, the proof of this is based on certain forbidden subgraph characterizations and Ramsey theoretic arguments, and the results are no longer meaningful if $\lambda$ grows faster than polylogarithmic in $n$. For highly structured graphs, such as strongly regular graphs (SRG), it is known [57] that if $|\lambda_{n}|$ is at most a small polynomial of the average degree, then the graph belongs to one of two special families. However, these results rely on the highly structured nature of SRGs. We refer the interested reader to the survey of Koolen, Cao, and Yang [56] for a general overview of the topic.
The smallest eigenvalue of the adjacency matrix also has theoretical importance. The celebrated Hoffman bound (see e.g. [43]) states that it controls the independence number of the graph. In particular, if $G$ is an $n$-vertex $d$-regular graph, then $\alpha(G)\leq\frac{n|\lambda_{n}|}{|\lambda_{n}|+d}$. Furthermore, the maximum of $|\lambda_{n}|$ and the second largest eigenvalue $\lambda_{2}$ determines the expansion and mixing properties of the graph [4], and as we will discuss later, $\lambda_{n}$ controls the maximum cut.
Many of these results show that the property of having small $|\lambda_{n}|$ and the existence of large trivial substructures, such as cliques, are interconnected. However, such results were previously only known when $|\lambda_{n}|$ is bounded by a constant, or growing very slowly with $n$. We prove that this phenomenon already starts to appear when $|\lambda_{n}|<n^{1/4-o(1)}$, and we show that graphs with smallest eigenvalue below this threshold converge to a trivial structure: a disjoint union of cliques. The exponent $1/4$ is also sharp, a celebrated construction of de Caen [26] related to equiangular lines provides a graph with smallest eigenvalue $|\lambda_{n}|=\Theta(n^{1/4})$ which is far from the disjoint union of cliques. We say that an $n$-vertex graph $G$ is $\mu$-close to some family of graphs $\mathcal{F}$ if the edit distance of $G$ to some member of $\mathcal{F}$ is at most $\mu n^{2}$.
@@ 1.3 Smallest eigenvalue
Let $\varepsilon,\delta>0$, then the following holds for every sufficiently large $n$. Let $G$ be an $n$-vertex graph such that $|\lambda_{n}|\leq n^{1/4-\varepsilon}$. Then $G$ is $\delta$-close to the disjoint union of cliques.
The previous theorem only ensures $o(1)$-closeness in case $|\lambda_{n}|\leq n^{1/4-o(1)}$. However, by requiring a slightly stronger upper bound on $|\lambda_{n}|$, we can also establish polynomial proximity to a disjoint union of cliques.
@@
Let $\varepsilon>0$, then there exists $\alpha>0$ such that the following holds for every sufficiently large $n$. Let $G$ be an $n$-vertex graph such that $|\lambda_{n}|\leq n^{1/6-\varepsilon}$. Then $G$ is $n^{-\alpha}$-close to the disjoint union of cliques.
While these results give strong structural results about somewhat dense graphs with small $|\lambda_{n}|$, they are no longer meaningful for sparse graphs $G$. On the other hand, it is not possible to formulate any reasonable extension of the previous theorems for sparse graphs, as the following example shows. The line graph of a graph always satisfies that $|\lambda_{n}|\leq 2$, but the line graph of the complete graph $K_{s}$ has $m=\Theta(s^{3})$ edges and it is not possible to add/remove $o(m)$ edges to get a disjoint union of cliques.
Despite this, we recall that Theorem 1.2 shows that large cliques, with size comparable to the average degree, do emerge in graphs of any sparsity and small $|\lambda_{n}|$. This suggests that such graphs might be close to the blow-up of much smaller graphs, and shows that trivial structures start to appear at any sparsity, assuming $|\lambda_{n}|$ is sufficiently small.
We prove Theorem 1.5 in Section 9, Theorem 1.6 in Section 13.1.
1.4 Alon–Boppana theorem
The Alon–Boppana theorem [63] is a cornerstone result of spectral graph theory. It states that if $G$ is an $n$-vertex $d$-regular graph, then the second largest eigenvalue $\lambda_{2}$ of the adjacency matrix is at least
$\lambda_{2}\geq 2\sqrt{d-1}-o_{n}(1).$
This result is often misquoted, with the $o_{n}(1)$ term forgotten or not understood properly. In its precise formulation, the Alon–Boppana theorem states that if $D$ is the diameter of $G$, then $\lambda_{2}\geq 2\sqrt{d-1}-\frac{2\sqrt{d-1}}{\lfloor D/2\rfloor}.$ In particular, if $D\rightarrow\infty$, which is satisfied in case $d=n^{o(1)}$, one gets the former lower bound. For fixed $d$, families of graphs satisfying $\max\{|\lambda_{n}|,\lambda_{2}\}\leq 2\sqrt{d-1}$ are called Ramanujan graphs, and their existence is known for many different values of $d$ [47]. A breakthrough of Friedman [38] shows that random $d$-regular graphs are close to being Ramanujan. Since the spectral gap $d-\lambda_{2}$ controls the expansion properties of the graphs, the Ramanujan graphs are optimal expanders. For this reason, such graphs are of great interest in the design of resilient networks, with countless further applications in theoretical computer science and extremal combinatorics.
In the case where the diameter $D$ is at most three, which can already happen if $d\approx n^{1/3}$, the Alon–Boppana bound is no longer meaningful. Also, one cannot hope for the bound $\lambda_{2}=\Omega(\sqrt{d})$ to hold unconditionally; for example the complete bipartite graph has $\lambda_{2}=0$. Recently, a number of authors [8, 11, 48, 64] studied the second eigenvalue in the case of denser graphs, and uncovered some highly unexpected behavior of its extremal value. In particular, [64] (see also [11] for a short note) proved that $\lambda_{2}=\Omega(d^{1/2})$ continues to hold for $d\leq n^{2/3}$, however, $\lambda_{2}=\Omega(n/d)$ for $d\in[n^{2/3},n^{3/4}]$, and this is (essentially) sharp by an old strongly regular graph construction of Metz (see [75]). Moreover, for $d\in[n^{3/4},(1/2-\varepsilon)n]$, we have $\lambda_{2}\geq\Omega_{\varepsilon}(d^{1/3})$, which is also sharp for $d=\Omega(n)$ by recent constructions of Davis, Huczynska, Johnson, and Polhill [25]. As we observed earlier, if $d=n/2$, we might have $\lambda_{2}=0$ by the complete bipartite graph. In general, when d=(11/r)nd=(1-1/r)n for some positive integer $r$, then the Turán graph $T_{r}(n)$, the complete $r$-partite graph with parts of size $n/r$, is $d$-regular and satisfies $\lambda_{2}=0$.
However, what happens when $d$ is not of the form (11/r)n(1-1/r)n, or $G$ is far from a Turán graph? The methods of [64] and related papers no longer apply when $d>n/2$, and there are no obvious further obstructions for having large second eigenvalue. In [64], it was conjectured that the answer to the second question is $\Omega(n^{1/4})$, which is then sharp by the equiangular lines construction of de Caen [26]. Considering complements, Theorem 1.5 immediately implies an almost complete solution of this conjecture. If $G$ is a regular graph with second eigenvalue $\lambda_{2}$, then the complement of $G$ has smallest eigenvalue $\lambda_{2}-1$.
Theorem 1.7.
Let $\varepsilon,\delta>0$, then the following holds for every sufficiently large $n$. Let $G$ be an $n$-vertex $d$-regular graph such that $\lambda_{2}\leq n^{1/4-\varepsilon}$. Then $G$ is $\delta$-close to a Turán graph. Thus, if $\lambda_{2}<n^{1/4-\varepsilon}$, then
@@ 1.4 Alon–Boppana theorem
1.5 Bisection width
$\operatorname{dfc}(G)=\begin{cases}\Omega(\sqrt{d}n)&\mbox{ if }d\leq n^{2/3},\\ \Omega(n^{2}/d)&\mbox{ if }d\in[n^{2/3},n^{4/5}],\\ \tilde{\Omega}(d^{1/4}n)&\mbox{ if }d\in[n^{4/5},(1/2-\varepsilon)n].\end{cases}$
These results are sharp for $d\in[1,n^{3/4}]$, and there are $d$-regular graphs for $d\approx n/3$ with deficit $O(n^{4/3})$. For d=n(11/r)d=n(1-1/r), where $r$ is a positive integer, the Turán graph $T_{r}(n)$ shows that we cannot hope for a bound better than $\Omega(n)$. Räty, Sudakov, and Tomon [64] conjectured that Turán graphs are the only obstruction to large deficit. Using the terminology of positive discrepancy, they conjectured that if $\operatorname{dfc}(G)=o(n^{5/4})$, then $G$ is $o(1)$-close to a Turán graph. We prove that this conjecture holds qualitatively, by establishing the bisection width analogue of our MaxCut result (Theorem 1.4).
Theorem 1.8.
The proof of Theorem 1.8 is presented in Section 13.2.
2 Proof overview
First, we outline the proof of Theorem 1.5, that is, that if a graph $G$ has smallest eigenvalue $|\lambda_{n}|\leq n^{1/4-\varepsilon}$, then $G$ is $\delta$-close to a disjoint union of cliques. Let $A$ be the adjacency matrix of $G$. We study the identity $A=A\circ A$ from a spectral perspective, where $\circ$ denotes the Hadamard product (see Section 4 for formal definitions). Writing $A=\sum_{i=1}^{n}\lambda_{i}v_{i}v_{i}^{T}$ for the spectral decomposition, we get that
@@ 2 Proof overview
But how to use this identity? An instructive case is when $G$ is a Cayley graph on some finite abelian group $(\Gamma,+)$. In this case, (1) reduces to a clean convolution relation: $\lambda_{a}=\frac{1}{n}\sum_{b+c=a}\lambda_{b}\lambda_{c}$ for all $a\in\Gamma$. Here, the eigenvalues are re-indexed by the group elements, and the identity follows from special properties of the characters of the group, which are also the eigenvectors of $G$. See Brouwer and Haemers [16] for further details. We can almost ignore the negative terms in the sum since the smallest eigenvalue is bounded. Hence, this convolution relation basically tells that large eigenvalues reinforce each other, i.e. if $\lambda_{b},\lambda_{c}\geq T$, then $\lambda_{b+c}\gtrsim T^{2}/n$. This motivates the definition $S_{T}=\sum_{\lambda_{i}\geq T}\lambda_{i}$, the spectral weight above threshold $T$. Summing over all $\lambda_{b}\geq T$ and $\lambda_{c}\geq T$, the above observation essentially gives that
@@
Surprisingly, this does generalize to arbitrary graphs to give the curious recursive inequality on the sum of large eigenvalues: for all $T\geq 2|\lambda_{n}|\sqrt{n}$,
@@
We achieve this by compressing both sides of (1) to the subspace $W$ spanned by the vectors $v_{i}\circ v_{j}$ where $\lambda_{i},\lambda_{j}\geq T$; see Section˜6.1 for a detailed argument. In terms of the quadratic sum of the eigenvalues, this implies that most contribution comes from large eigenvalues; we show this in Section˜6.3. But this means that $A$ is close to a low-rank positive semidefinite matrix in the Frobenius norm, which is only possible if $G$ is close to a disjoint union of cliques. We prove this in Section˜9.
Now we discuss the proof of Theorem 1.6, which requires a lot more work. Recall that this theorem states that if a graph $G$ has smallest eigenvalue $|\lambda_{n}|\leq n^{1/6-\varepsilon}$, then $G$ is $n^{-\alpha}$-close to a disjoint union of cliques. The bottleneck in the previous argument is the last part, it requires that the rank of the low-rank approximation of $A$ is at most constant, in which case we can establish $o(1)$-closeness to a disjoint union of cliques. In order to overcome this, we first aim to show that either $G$ is already sparse (in which case $G$ is $n^{-\alpha}$-close to the empty graph), or $G$ contains very large cliques. If we are able to do this, then we repeatedly pull out large cliques, which gives enough structure to easily conclude the desired result. In order to find our large cliques, we go through 4 phases of densification, i.e. we find denser and denser subgraphs of $G$. We keep in mind that the Cauchy interlacing theorem ensures that the induced subgraphs of $G$ also have large smallest eigenvalue. In what follows, we use $c$ to denote a small positive constant depending only on $\varepsilon$ and $\alpha$, and different occurrences of $c$ might denote different quantities.
@@
We show that $G$ contains an unusually high number of triangles. We count triangles by the cubic sum of eigenvalues, and we argue that this sum is large because most of the mass of the quadratic sum of eigenvalues is concentrated on the few largest eigenvalues. Having many triangles means that we can find a vertex whose neighbourhood is much denser than $G$. We repeat this process until we find a subgraph $G_{1}$ on $n^{1-c}$ vertices of positive constant density. This phase of the argument requires $|\lambda_{n}|\leq n^{1/6-\varepsilon}$. This can be found in Section 8.
@@
In Theorem 1.5, we already established that $G_{1}$ is $o(1)$-close to a disjoint union of cliques. Therefore, if $G_{1}$ has positive constant edge density, this implies that $G_{1}$ contains a linear size subgraph $G_{2}$ of edge density 1o(1)1-o(1). This can be found in Section 9.
@@
For very dense graph, we employ a new method, inspired by the works of [64]. Let $G$ be an $n$-vertex graph of density 1p1-p, with $10^{-5}>p>n^{-1/2}$, and $|\lambda_{n}|\leq n^{1/6-o(1)}$. We consider the matrix
$B=A-\lambda_{1}v_{1}v_{1}^{T}+|\lambda_{n}|I,$
and we study the triple Hadamard product $B\circ B\circ B$. The matrix $B$ is positive semidefinite, so the Schur product theorem ensures that $B\circ B\circ B$ is also positive semidefinite. But then $x^{T}(B\circ B\circ B)x\geq 0$ for every vector $x$. We choose $x$ to be the characteristic vector of a carefully chosen, linear sized set $U\subset V(G)$, and argue that the previous inequality can be only satisfied if the density of $G[U]$ is at least $1-O(p^{3})$. Therefore, repeating this argument on $G_{2}$, we get a rapid density increment until we reach a graph $G_{3}$ of edge density $1-n^{-1/2}$. This can be found in Section˜10.
Phase 4.
Above edge density $1-n^{-1/2}$, we may consider simple identities involving the eigenvalues. Let $G$ be a graph of edge density at least $1-n^{-1/2}$ and $|\lambda_{n}|\leq n^{1/6}$, and consider the four identities
$\Delta(G)\geq\lambda_{1},\ \ \ \ \sum_{i=1}^{n}\lambda_{i}=0,\ \ \ \ \sum_{i=1}^{n}\lambda_{i}^{2}=2e(G),\ \ \ \ \sum_{i=1}^{n}\lambda_{i}^{3}\geq 0.$
We find an almost regular subgraph $G_{4}$ of $G_{3}$, and use simple algebraic manipulations to argue that the expressions above can be only satisfied if $G_{4}$ has density at least $1-n^{-1+c}$. At this point, we simply apply Turán’s theorem to find a clique of size $n^{1-c}$. This can be found in Sections˜11 and 12.
Next, we discuss Theorem 1.2, which states that any graph with average degree $d$ and smallest eigenvalue $|\lambda_{n}|\leq d^{\gamma}$ contains a clique of size $d^{1-O(\gamma)}$. In order to prove this, we introduce a phase 0 of the densification process, which immediately lets us move to density at least $n^{-c}$, assuming $|\lambda_{n}|$ is sufficiently small with respect to the average degree. Then, we apply the previous four densification steps to conclude the proof.
@@
Let $G$ be a graph of average degree $d$, then we show that $G$ contains a subgraph on $d$ vertices of edge density $\Omega(1/|\lambda_{n}|)$. This follows by picking a vertex $x$ with a set of $d$ neighbours $S$, and then analyzing $v^{T}Av$ for an appropriately chosen $v$ with support $\{x\}\cup S$. This can be found in Section˜7.
In order to prove our results concerning graphs with small maximum cut, that is, Theorems 1.3 and 1.4, we follow the same steps. In Section 5, we present a toolkit that gives various lower bounds on the MaxCut based on the negative eigenvalues of the graph. With the help of these, instead of having a bound on $|\lambda_{n}|$, we can bound the sum, quadratic sum, and cubic sum of the negative eigenvalues. This allows us to transfer most of the machinery developed for graphs with bounded smallest eigenvalue to graphs with bounded MaxCut, but with the cost of incurring some losses quantitatively.

### v2

@@ 1.3 Smallest eigenvalue
Let $\gamma\in(0,1/4)$, $\delta>0$, and let $n$ be sufficiently large. If $G$ is an $n$-vertex graph with $|\lambda_{n}|\leq n^{\gamma}$, then $G$ is $\delta$-close to the vertex-disjoint union of cliques.
The previous theorem only guarantees $o(1)$-closeness in the case $|\lambda_{n}|\leq n^{1/4-o(1)}$. Note that one cannot expect a substantially stronger result than $o(1)$-closeness, as illustrated by the following example. Suppose $G$ is the disjoint union of $n^{4\varepsilon}$ copies of the graph constructed by de Caen [30], each of size $n^{1-4\varepsilon}$. Then $G$ satisfies $|\lambda_{n}|\leq n^{1/4-\varepsilon}$ and, as mentioned above, one must add or remove $\Omega(n^{4\varepsilon}(n^{1-4\varepsilon})^{2})$ edges to make $G$ a disjoint union of cliques. Therefore, one cannot hope to prove that $G$ is closer than $n^{-4\varepsilon}$ to a union of cliques. By imposing a slightly stronger bound on $|\lambda_{n}|$, we can indeed establish polynomial proximity to a disjoint union of cliques.
@@
For every $\gamma\in(0,1/6)$, there exists $\alpha>0$ such that for every sufficiently large $n$ we have the following. If $G$ is an $n$-vertex graph with $|\lambda_{n}|\leq n^{\gamma}$, then $G$ is $n^{-\alpha}$-close to the vertex-disjoint union of cliques.
While these results provide good structural understanding of somewhat dense graphs with small $|\lambda_{n}|$, they are no longer meaningful for sparse graphs. In fact, it is impossible to formulate any reasonable extension of the previous theorem even for moderately sparse graphs, as the following example show. All line graphs satisfy $|\lambda_{n}|\leq 2$, but the line graph of the complete graph $K_{t}$ has $\Theta(t^{2})$ vertices and $m=\Theta(t^{3})$ edges and it is not possible to add or remove $o(m)$ edges to get a disjoint union of cliques.
Despite this, we recall that Theorem˜1.2 shows that large cliques, of size $d/|\lambda_{n}|^{4}$, emerge in graphs of any sparsity, whenever $|\lambda_{n}|\leq d^{\gamma}$ for $\gamma\in(0,1/10)$. This extends the result of Yang and Koolen [86], who showed that if $d$ is exponentially large compared to $|\lambda_{n}|$, then $G$ must contain a clique of size $d/|\lambda_{n}|^{3}$. This suggests that such graphs might be close to the blow-up of much smaller graphs, and shows that trivial structures start to appear at any sparsity, assuming $|\lambda_{n}|$ is sufficiently small. We prove Theorem˜1.4 in Section˜8 and Theorem˜1.5 in Section˜11.
1.4 Alon–Boppana theorem
The Alon–Boppana theorem [73] is a cornerstone result of spectral graph theory. It states that if $G$ is an $n$-vertex $d$-regular graph, then the second-largest eigenvalue $\lambda_{2}$ of the adjacency matrix is at least
$\lambda_{2}\geq 2\sqrt{d-1}-o_{n}(1).$
In its precise formulation, the Alon–Boppana theorem states that if $D$ is the diameter of $G$, then $\lambda_{2}\geq 2\sqrt{d-1}-\frac{2\sqrt{d-1}}{\lfloor D/2\rfloor}.$ In particular, if $D\rightarrow\infty$, which is satisfied in the case $d=n^{o(1)}$, one gets the former lower bound. For fixed $d$, families of graphs satisfying $\max\{|\lambda_{n}|,\lambda_{2}\}\leq 2\sqrt{d-1}$ are called Ramanujan graphs, and their existence is known for many different values of $d$ [55]. A breakthrough of Friedman [44] shows that random $d$-regular graphs are close to being Ramanujan. Since the spectral gap $d-\lambda_{2}$ controls the expansion properties of graphs, the Ramanujan graphs are optimal expanders. For this reason, such graphs are of great interest in the design of resilient networks, with countless further applications in theoretical computer science and extremal combinatorics.
In the case where the diameter $D$ is at most three, which can already happen if $d\approx n^{1/3}$, the Alon–Boppana bound is no longer meaningful. Also, one cannot hope for the bound $\lambda_{2}=\Omega(\sqrt{d})$ to hold unconditionally; for example the complete bipartite graph has $\lambda_{2}=0$. Recently, a number of authors [10, 13, 56, 74] studied the second eigenvalue in the case of denser graphs, and uncovered some highly unexpected behavior of its extremal value. In particular, [10, 74] (see [13] for a short note) proved that
$\lambda_{2}=\begin{cases}\Omega(d^{1/2})&\mbox{ if }d\leq n^{2/3},\\ \Omega(n/d)&\mbox{ if }d\in[n^{2/3},n^{3/4}],\\ \Omega(d^{1/3})&\mbox{ if }d\in[n^{3/4},(1/2-\varepsilon)n].\end{cases}$
Moreover, these bounds are (essentially) sharp in the first two regimes, and also in case $d=\Omega(n)$ [29]. As we observed earlier, if $d=n/2$, we might have $\lambda_{2}=0$ by the complete bipartite graph. In general, when d=(11/r)nd=(1-1/r)n for some positive integer $r$, the Turán graph $T_{r}(n)$ (the complete $r$-partite graph with parts of size $n/r$) is $d$-regular and satisfies $\lambda_{2}=0$.
However, what happens when $d$ is not of the form (11/r)n(1-1/r)n or $G$ is far from a Turán graph? The methods of [74] and related papers no longer apply when $d>n/2$, and there are no obvious further obstructions for having large second eigenvalue. In [74], it was conjectured that the answer to the second question is $\Omega(n^{1/4})$, which is sharp by the equiangular lines construction of de Caen [30]. Considering complements, Theorem 1.4 immediately implies an almost complete solution of this conjecture. If $G$ is a regular graph with second eigenvalue $\lambda_{2}$, then the complement of $G$ has smallest eigenvalue $\lambda_{2}-1$.
@@ 1.4 Alon–Boppana theorem
Let $\gamma\in(0,1/4)$, $\delta>0$, and let $n$ be sufficiently large with respect to $\gamma,\delta$. If $G$ is an $n$-vertex $d$-regular graph with $\lambda_{2}\leq n^{\gamma}$, then $G$ is $\delta$-close to a Turán graph. Thus, if $\lambda_{2}\leq n^{\gamma}$, then
@@
2 Proof overview and organization
First, we outline the proof of Theorem˜1.4, which states that if a graph $G$ has smallest eigenvalue $|\lambda_{n}|\leq n^{\gamma}$ for some $\gamma\in(0,1/4)$, then $G$ is $\delta$-close to a disjoint union of cliques. Let $A$ be the adjacency matrix of $G$. In order to exploit the fact that $A$ is a $0/1$-matrix, we study the identity $A=A\circ A$, where $\circ$ denotes the entry-wise or Hadamard product (so that $(A\circ B)_{ij}=A_{ij}B_{ij}$, see Section 4 for formal definitions). Writing $A=\sum_{i=1}^{n}\lambda_{i}v_{i}v_{i}^{T}$ for the spectral decomposition, we get that
@@ 2 Proof overview and organization
But how to use this identity? An instructive case is when $G$ is a Cayley graph of a finite abelian group $(\Gamma,+)$. In this case, the eigenvalues can be indexed by the group elements, and (1) reduces to a clean convolution relation: $\lambda_{a}=\frac{1}{n}\sum_{b+c=a}\lambda_{b}\lambda_{c}$ for all $a\in\Gamma$. This identity also follows from special properties of the characters of the group, which are also the eigenvectors of $G$ (see Section 1.4.9 of Brouwer and Haemers [19] for further details). If $\lambda_{n}$ is not very negative, we can almost ignore the negative terms in the sum. Hence, roughly speaking, this convolution relation shows that large eigenvalues reinforce each other, i.e. if $\lambda_{b},\lambda_{c}\geq T$, then $\lambda_{b+c}\gtrsim T^{2}/n$. This motivates the definition $S_{T}=\sum_{\lambda_{i}\geq T}\lambda_{i}$, the spectral weight above threshold $T$. Summing over all $\lambda_{b}\geq T$ and $\lambda_{c}\geq T$, the above observation gives that
@@
This heuristic can be converted into a formal argument and generalized to arbitrary graphs, yielding the following curious recursive inequality on the sum of large eigenvalues: for all $T\geq 2|\lambda_{n}|\sqrt{n}$,
@@
To derive (2) in general graphs, we use the notion of subspace-compression of matrices. We compress both sides of (1) onto the subspace $W$ spanned by the vectors $v_{i}\circ v_{j}$ where $\lambda_{i},\lambda_{j}\geq T$; see Section˜6.1 for a detailed argument. We then use the recursive inequality (2) to show that the contribution of small eigenvalues in the quadratic sum of all eigenvalues is negligible; we show this in Section˜6.3. But this means that $A$ can be well-approximated in Frobenius norm by a low-rank positive semidefinite matrix. However, this is only possible if $G$ is close to a disjoint union of cliques, which we prove in Section˜8.
Now we discuss the proof of Theorem 1.5, which states that if a graph $G$ has $|\lambda_{n}|\leq n^{\gamma}$ for some $\gamma\in(0,1/6)$, then $G$ is $n^{-\alpha}$-close to a disjoint union of cliques. The bottleneck in the previous argument is its last step, where we show that if $A$ is well-approximated by a low-rank matrix, then $G$ must be close to a union of cliques. Our argument requires the rank of the approximation to be constant, which we cannot achieve if the graph $G$ is sparse. To overcome this, we first show that either $G$ is already $n^{-\alpha}$-sparse (in which case $G$ is $n^{-\alpha}$-close to the empty graph), or $G$ contains a very large clique. We then repeatedly pull out large cliques, which gives enough structure to easily conclude the desired result. In order to find large cliques, we use a density-increment strategy, which is divided into three phases. We use $\varepsilon$ to denote a small positive constant depending only on $\gamma$ and $\alpha$.
@@
Using (2), we show that $G$ contains an unusually high number of triangles. We count triangles by the cubic sum of eigenvalues, and we argue that this sum is large because most of the mass of the quadratic sum of eigenvalues is concentrated on the few largest eigenvalues. Having many triangles means that we can find a vertex whose neighbourhood is much denser than $G$. We repeat this process until we find an induced subgraph $G_{1}\subseteq G$ on $n^{1-\varepsilon}$ vertices of positive constant density. This phase of the argument requires $\gamma<1/6$. The details are given in Section 7.
@@
Due to Cauchy’s interlacing theorem, $G_{1}$ also lacks very negative eigenvalues. Hence, Theorem 1.4 applies to $G_{1}$, implying that $G_{1}$ is $o(1)$-close to a disjoint union of cliques. Therefore, using that $G_{1}$ has positive constant edge density, we show that $G_{1}$ contains a linear-sized induced subgraph $G_{2}\subseteq G_{1}$ of edge density 1o(1)1-o(1). This step is explained in Section 8.
@@
For very dense graphs, we employ a new method, inspired by the work of Räty, Sudakov and Tomon [74]. We prove that if $G_{3}$ is a somewhat regular induced subgraph of $G_{2}$, then the complement of $G_{3}$ must have average degree $O(|\lambda_{n}|^{2})$, assuming $|\lambda_{n}|\ll|V(G_{3})|^{1/4}$. Thus by Turán’s theorem, $G_{3}$ contains a clique of size $\Omega(|V(G_{3})|/|\lambda_{n}|^{2})$. In order to prove this, we analyse the triple Hadamard product of certain positive semidefinite shifts of the adjacency matrix. This can be found in Section 9.
We put together all of these ingredients in Section 11 to provide the proof of Theorem 1.2.
Next, we discuss Theorem˜1.2, which states that any graph with average degree $d$ and smallest eigenvalue $|\lambda_{n}|\leq d^{\gamma}$ contains a clique of size $d^{1-O(\gamma)}$. Note that the methods discussed above only apply to somewhat dense graphs, whose average degree is $n^{1-\alpha}$ for some small $\alpha$. Therefore, as the first step in proving Theorem˜1.2, we introduce another densification method, allowing us to move to density at least $1/|\lambda_{n}|$, which will be sufficient assuming $\lambda_{n}$ is small with respect to the average degree. Then, we apply the previous three densification steps to conclude the proof. More precisely, we prove the following.
@@
If $G$ has average degree $d$, then we show that $G$ contains a subgraph on $d$ vertices of edge density $\Omega(1/|\lambda_{n}|)$. This follows by picking a vertex $x$ with a set of $d$ neighbours $S$, and then analyzing the inequality $v^{T}Av\geq\lambda_{n}\|v\|_{2}^{2}$ for an appropriately chosen $v$ with support $\{x\}\cup S$. This can be found in Section˜10.
In order to prove our results concerning graphs with small maximum cut, that is, Theorem 1.3, we follow the same steps. In Section 5, we present a toolkit that gives various lower bounds on MaxCut based on the negative eigenvalues of the graph. With the help of these, instead of having a bound on $|\lambda_{n}|$, we can bound the linear, quadratic, and cubic sum of the negative eigenvalues. This allows us to transfer most of the machinery developed for graphs with bounded smallest eigenvalue to graphs with bounded MaxCut, but with the cost of incurring some losses quantitatively.